A survey done for a magazine found that 77% of bike riders sometimes ride without a helmet. If four bike riders are randomly selected, find the probability that at least one of the riders does not wear a helmet at all times.
Answer by stanbon(75887)

A survey done for a magazine found that 77% of bike riders sometimes ride without a helmet. If four bike riders are randomly selected, find the probability that at least one of the riders does not wear a helmet at all times.
------
Binomial Problem with n = 4 and p(helmet) = 0.77 ; p(no helmet) = 0.23
-----
P(x >=1 no helmet) = 1 - P(all wear helmets) = 1- 0.77^4 = 0.6485
===================
Cheers,
Stan H.
===================